The perimeter of an isosceles triangle is 100 centimeters. If two sides are equal length and the third side is 10 centimeters longer than the others, find the length of the three sides
Answer by checkley71(8403)

100=(x+10)+x+x or 3x+10=100 or 3x=100-10 or 3x=90 or x=90/3 or x=30
proof 100=(30+10)+30=30 or 100=40+30+30 or 100=100